OpenStudy (anonymous):

tan (1/2x) = t, then sin x =??

OpenStudy (anonymous):

by \(\tan x =\frac{2t}{1-t^2}\), where t=tan(x/2)  so \(\sin x =\frac{2t}{1+t^2}\)